News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 13 September 2002 9.00am EDT

During the last week the observing schedule was interrupted following a high radiation event at 1:49pm EDT on Sep 7 that stopped the daily load and activated the science instrument safing sequence (SCS 107). All actions were nominal and radiation levels dropped sufficiently to allow the loads to be restarted at 4:23pm on Sep 9. Observations of PSS0808+5215, PSS 1507+5603, AX J1531.8+2414 and Q2127+040 (total 30.6 Ks) were impacted by the event and will be rescheduled in a later load.

Chandra passed through the lunar eclipse on Sept 7 with nominal power and thermal performance. Of note was that the radiation safing ocurred only 15 minutes after the eclipse exit, however appropriate contingencies were in place to handle a trip at any time through the eclipse. Real-time procedures were completed on Sept 8 to back-out the eclipse and contingency software patches and clear a glitch counter that incremented during eclipse exit.

A Chandra image release of the Tycho supernova remnant was made on Sept 6. For details see: http://chandra.harvard.edu/photo/2002/0005/index.html.

The observing schedule for next week is shown below.

------------------------------------------
Radiation Belts                      Sep 15
AX J1511.7+5702         ACIS-S       Sep 16
PSS 1342+4944           ACIS-S
A2259                   ACIS-I
GRS 1758-258            HRC-I
PSR J1420-6048          ACIS-S
PSR J1420-6048          ACIS-S
MACSJ0152.5-2852        ACIS-I       Sep 17
Pictor A - West         ACIS-S
NGC 1332                ACIS-S
Radiation Belts                      Sep 18
NGC1600                 ACIS-S
AX J0844.8+5004         ACIS-S       Sep 19
NGC1313 SA and SB       HRC-I
NGC 1332                ACIS-S
NGC1600                 ACIS-S
NGC2146                 ACIS-S       Sep 20
3C401                   ACIS-S
H230236                 ACIS-I
Radiation Belts
3C401                   ACIS-S       Sep 21
NGC7052                 ACIS-S
PSR J1420-6048          ACIS-S
DEM L152                ACIS-S       Sep 22
BR B0305-4957           ACIS-S
3CR184 Cluster          ACIS-S
BR0308-1734             ACIS-S
Pictor A - West         ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports